I need help with identifying the unknown acid, percent error of pka1, and pka2

I need help with identifying the unknown acid, per
Answer #1

Your data matched with Arsenic acid pKa's

the difference in pKa1 = 2.26-2.85 = -0.59

the percent error for pKa1 = 26% lower then expected value

the difference in pKa2 = 6.77-6.15 = 0.62

The percent error for pKa2 = 9.15% higher then expected value

a. Methyl orange has pH range of indication 3.1-4.4 So it would only show color change at low pH

b. Phenolphthalein has pH range 8.0-10.0 so it would show color change and can be used as indicator at higher pH ranges only say 8.25 as in the above case.

c. Lastly mlachite green has a color indication range pH 0.2-1.8, so it can also be used at much low pH values.

Looking at the above data we see, pheolphthalein works as expected changes color from colorless to pink at 8.25 pH

Methyl orange also works as expected shows color change arounf pH 4

malachite green on the other hand does not work in either case since its pH range is very different from the equivalence points and thus the pH of the acid used for titration.